# Deformable Cluster Manipulation via Whole-Arm Policy Learning

### v1

@@ III-F Observation & Reward Space
Proprioceptive joint pose: ${q}$ 6
@@
Our observation space $O$ consists of five feature sets, listed in Table I. First, we construct two distinct feature sets from point cloud information: the KME features described in Section III-D, which characterise global structure, and a local feature group, which focuses on nearby point distances. The latter includes the Euclidean distance of the end-effector to the closest $k=5$ branch points and a safety breach indicator to reduce the robot-line contacts, computed as:
@@
As the branches overhanging the power line already act as conductors [14], the robot touching the power line in itself does not compromise the validity of our approach; nevertheless, a safety margin $d_{sm}=4$cm is introduced to reduce the wear and tear to the line through reward shaping.
A novel aspect of our method is its multi-modality, attributable to a whole-arm touch detector. Recently, [11] has shown that in the absence of expensive tactile sensors, time-series sliding window torque measurements: $[\tau_{t},\tau_{t-1},\tau_{t-2},...,\tau_{t-(m-1)}]$ over the last $m$ time steps along with other proprioceptive features, can be used to build a classifier trained to detect contact ‘bumps’ at the current time-step $t$, from labeled touch data. While [11] is trained with touch features for a simple reach task, we extend that approach and demonstrate that the touch classifier is useful even in the presence of vision and can aid complex manipulation tasks. During simulation, a proxy classifier $\mathbb{I}(\|{F_{t}}\|_{2}>f_{u})$ computes the force threshold $f_{u}$ breaches by the robot links $l$, leveraging the net contact force $F_{t}\in\mathbb{R}^{l\times 3}$ exposed by Isaac Gym. The full body contact awareness from this touch detector is crucial for the arm to leverage its multiple links (e.g. forearm and wrist) for manipulation, to supplement the partial, noisy, and time-varying point clouds.
Our simple reward structure listed below has three components: a) a clearance reward $r_{h}$ to encourage manipulator towards lower occlusion scores $h_{t}$, b) a smoothness reward $r_{q}$ on the joint velocities $\dot{q}$, and c) a bonus for the time-steps when the arm links are clear of the power line safety margin,
@@
III-G Domain Randomisation
This work aims to generalise a learned manipulation policy to deformable scenarios with unseen geometry, undetermined dynamics, and devoid of chromatic affinity, by not relying on digital twins [27], parameter inference [29][4], and RGB images respectively during the training phase. Instead, we perturb simulation parameters, for example, L-system structural traits such as branch divergence angle and segment elongation rate, and dynamics parameters such as spring stiffness, damping and friction, among others. In particular, we sub-sample and permute all training point clouds while adding Gaussian noise during each time-step.
@@ III-G Domain Randomisation
Figure 3: Terminal state poses: (a)(b) Simulation trajectory states showing the whole arm being utilised to shield the power line. (c)(d) & (e) Similar real strategies using various arm links for clearance.
@@ IV-A Simulation Setting
For the power line clearance scenario illustrated in Fig. 2(d), we run 6144 Gym environments in parallel, each of which contains a tree with deformable branches (in green), a rigid power line (in red), and a Kinova arm actuated via velocity control of its six link joints, omitting the end-effector. Each tree is formed by drawing from a Gaussian distribution ($\sigma=0.1$) representing the L-system morphology parameters to ensure high diversity in topology but sufficient physical feasibility, in addition to arbitrary trunk rotation to expose the agent to varying branch occlusion patterns. Furthermore, the power line location is randomised, but constrained to the reach region of the gripper and semi-parallel (within ±$\pm 15^{\circ}$) to the axis connecting the trunk and the robot base. Finally, we perform all experiments with an average $h_{t}\in[0,1]$ of at least $0.7$ across all environments, noting that occlusion score can increase as well from the arm pushing more branches closer to the power line.
@@
A critical challenge within this framework is the contact explosion resulting from the combinatorial interactions among the multi-link arm, multi-branch tree, and the power line, exacerbated by a large number of training environments. This issue, common in contact-rich parallel simulations, can quickly overwhelm the simulation constraint solvers, triggering object inter-penetrations and CUDA memory overflows. It has been extensively studied by [37] and [38], which suggest complex CUDA-level penetration checkers and mesh-level contact binning, respectively. In contrast, we take a shortcut to contact reduction at the cost of a marginal drop in test performance. Specifically, during training, we allow the branches (but not the robot) to penetrate the power line through collision masking, thereby removing one key set of interactions. On the other hand, during simulation tests, all contacts are in place to accurately reflect real-world, but the environment count is set to a low value, 512 per test dataset. The technique mentioned above to allow for slight differences in training and test environments, along with careful tuning of the Temporal Gauss-Seidel (TGS) contact solver parameters, enables us to run large-scale training required for RL without compromising representational accuracy.
TABLE II: Feature Ablation
@@
Description $d_{\text{max}}$ Train Rew Trials Test Rew Test SROcc DropTouchSteps in Succ
(a) Proprioceptive + Heuristic $h_{t}$ 1495.6 1536 1413.6 ± 28.7 36.5 ± 2.3 18.0 ± 1.3 2.8 ± 0.3 237.7 ± 19.2
(b) (a) + Local PC Features 0.02 1531.7 1536 1486.5 ± 49.1 43.8 ± 4.5 25.2 ± 3.6 2.1 ± 0.4 258.3 ± 32.1
(c) (b) + Touch Feature 0.02 1606.6 1536 1506.1 ± 28.5 46.1 ± 2.6 27.9 ± 4.8 2.6 ± 0.8 268.9 ± 17.7
(d) (b) + KME Features 0.02 1739.1 1536 1630.2 ± 31.6 55.7 ± 1.1 40.1 ± 0.4 1.9 ± 0.1 348.2 ± 22.7
(e) All Features 0.02 1704.3 1536 1655.9 ± 26.4 63.4 ± 2.8 46.3 ± 3.3 3.8 ± 0.3 363.4 ± 20.6
(f) Final Model 0.005 1825.1 1536 1763.6 ± 21.5 68.4 ± 2.1 53.7 ± 2.2 3.3 ± 0.4 429.8 ± 25.6
@@ IV-B Hardware Design & Sim-to-Real
In this work, we aimed to train a policy in simulation, running on an NVIDIA RTX 4090, and transfer it to real without a policy adaptation phase. During inference, the agent interacts with our real Kinova 6-DOF Jaco 2 through the Kinova ROS API and a custom REST interface, running on a separate low-grade workstation at a 60Hz control rate. Unlike prior works focused on artificial plants with soft leafy resistance, our test branches are randomly chosen from real trees, ensuring sufficient stiffness to resist clearance. A single fixed Intel RealSense D405 camera, operating at 30 fps, provides the RGB-D ($480\times 640$) observation of the workspace, illustrated in Fig. 2(c). The segmented real branch point clouds are constructed from the RGB-D camera images; in contrast, simulation tree points are randomly sampled from the coarse-grained branch segment surfaces. The real robot point cloud is built once from meshes, just as in the simulation, and updated at each step with the 3D coordinate transformation tree from ROS.
Furthermore, to address the significant sim-to-real discrepancies in robot parameters, contact dynamics, and gravity compensation, exacerbated by velocity control, we leverage Segmented Steady State Error Control (SSED)[11]. This scheme alternates between applying an action $a_{t}\sim\pi_{w}(\cdot|s_{t})$ to the desired state $s^{d}_{t}$, instead of the current, but periodically synchronising $s^{d}_{t}$ to $s_{t}$, in effect, rejecting short-term steady-state errors and clearing long-term accumulated offsets.
@@ IV-C Real Vision Pipeline
This section outlines the key transformation steps required to convert the real-time streaming images into RL features. First, we use Grounding DINO [39] to locate the objects in the scene with simple text prompts. The resulting bounding boxes are passed to the Segment Anything SAM-HQ [40] model to generate object masks zero-shot. Segmented masks are then de-projected to the robot base frame, leveraging the depth channel input. A key concern at this stage is that manipulators operating in velocity control mode have a control rate lower bound, 60Hz in our case, required to maintain velocity. Current state-of-the-art grounding and segmentation tools are unable to achieve this without significantly compromising the mask quality. Therefore, we use a video object segmentation framework Cutie [41], to propagate the mask across time-steps, occasionally replacing it with updated masks to prevent quality loss while tracking. Moreover, we use Python Shared Memory modules and asynchronous queues to go beyond the camera-constrained fps and meet the control frequency requirement. Notably, such a pipeline is problem agnostic; specifically, the grounding text prompt alone needs to be modified to switch from the power line context to a cable inspection scenario, a promising direction for future work.
@@ V EXPERIMENTS AND RESULTS
Experiment 1: First, we perform a simulation ablation study to justify our feature choices with the setup from Section IV-A. For each combination, we train independent policies by injecting varying levels of Gaussian Noise $\mathcal{N}(0,\sigma^{2})$ into the point clouds with the $3\sigma$ rule, i.e., $\sigma=d_{\text{max}}/3$, $d_{\text{max}}$ being the maximum deviation representative of the camera depth inaccuracy. The ablation results are shown in Fig. 4 and Table II. The box plot data points are policies constructed with $d_{\text{max}}$ levels ranging from $0.005$ to $0.04$, displaying the distribution of test rewards and success rates on a single unseen test set. On the other hand, each row in Table II is an individual policy tested on multiple test sets, listing a variety of additional evaluation metrics, with rows (a)-(e) showing the results applying a median noise $d_{\text{max}}=0.02$, representative of the ablation study. By contrast, row (f) shows our best model with a lower $d_{\text{max}}=0.005$, the depth inaccuracy of the D405 camera used for real experiments.
@@
The tabulated evaluation metrics are described as follows: Training Reward (Train Rew) and Test Reward (Test Rew) assess the agent’s performance during training and on unseen test scenarios, respectively. Test Success Rate (Test SR %) quantifies the percentage of test environments (Trials) where all occlusions were removed from the power line. Given that the branch collection will continue to resist clearance even after success is reached, we choose a conservative Test SR, considering a task as a success only if the line is occlusion-free for at least 10 consecutive steps. Occlusion reduction (Occ Drop %) shows the percentage drop in heuristic ($h_{t}$) from the start to end of test episodes averaged across all environments, while the (Touch %) reflects the frequency of undesirable instances where the arm breaches the pre-set safety margin $\delta_{sm}$ around the power line. The final metric (Steps in Succ) measures the efficiency of successful trajectories; as an example, the final row (f) implies that, on average, in a $1000$ step trajectory, the arm starting from the home position, traversed to the occlusion locations and moved the occluding branches away from the line in $1000-429.8=570.2$ steps, while the remaining $429.8$ steps were completely occlusion free for potential pruning, despite continued cluster resistance. All metrics except the power line touch (Touch %) are higher, the better.
@@
The plot in Fig. 4 demonstrates that both the KME feature group capturing the global point cloud attributes and the contact detection classifier form essential constituents of our multi-modal policy, and removal of either would hurt performance. Along the same lines, Table. II shows that despite the stochasticity of individual policies and high noise injection, the progressive addition of feature groups induces improvement across all evaluation metrics. To highlight, we report all test success rates without distinguishing between feasible and infeasible tasks, i.e., all occlusion patterns are randomly generated and may contain scenarios that cannot be solved by a single arm. In some unsolvable instances that contribute to the failures, the arm cannot generate sufficient torques (due to joint limits) to overcome the collective resistance of obstructing branches; in others, multiple disconnected branches can occlude the line at locations not closely spaced, only one of which can be removed at a time.
@@
Experiment 2: Second, we evaluate the representational efficiency of our approach to other point cloud feature extractors, namely, a Pytorch variant [42] of the well-known PointNet++ [35] and the Point2Vec [43]. Specifically, we replace the four global point features from Table I, i.e., $\overline{\mu_{\mathbb{P}}}[P^{wbr}]$, $\overline{\mu_{\mathbb{P}}}[P^{zbr}]$ , $\overline{\mu_{\mathbb{P}}}[P^{clr}]$ and $\overline{\mu_{\mathbb{P}}}[P^{rob}]$ with embeddings built from PointNet++ and Point2Vec. In each case, we use pre-trained checkpoints, down-sample the input points, and employ PCA post-feature extraction wherever necessary for consistency with our approach. We train each set for 250 epochs with 1024 environments, and the corresponding time taken (Train Time) and the accumulated training reward (Train Rew) at the end are tabulated. In addition, we compute ${GF_{t}^{(E)}}$ denoting the average time taken to build the 4x global point features only for a single RL time-step for $E$ environments. $GF_{t}^{(1024)}$ is indicative of the representation efficiency while training on 1024 environments and $GF_{t}^{(1)}$ represents the inference performance for one environment. The results are tabulated in Table IV, where lower is better for the time metrics, such as, $GF_{t}^{(1024)}$, $GF_{t}^{(1)}$, and Train Time while higher values are better for others.
@@
Comparing the time taken for a single time-step and for the overall training, from Table IV, our kernel-based approach shows orders of magnitude improvement, with similar training rewards. For the PointNet++ and Point2Vec cases, the maximum possible environments (Max Envs) we could spin up in our hardware was 1024, an indication of the GPU memory constraint due to the heavy pre-trained checkpoints. Notice that our final KME-based model training, row (f) from Table II, which uses 6x gym environments and 5x iterations compared to this experiment, takes just 4.5 hours to complete on an RTX 4090. Furthermore, the tabulated $GF_{t}^{(1)}$ values indicate that both PointNet++ and Point2Vec cannot meet the 60Hz real-time inference frequency requirement, justifying our approach. However, we acknowledge that improved efficiency may be at the cost of expressivity and that advanced baseline versions or quantisation techniques can reduce this time difference.
@@
Experiment 3: Finally, we go beyond simulation validation to evaluate our approach in a real-world setting under laboratory conditions. In real, we test two branches from different species, modifying either the branch orientation or the power line position in each experiment. Furthermore, in each test, we start with an $h_{t}$ of at least $0.8$, and consider both partial occlusion clearance and any touch to the power line as failure instances. The results are in Table IV.
@@
Corresponding real videos (see supplement and Fig. 3) demonstrate that the simulation-trained policy can adapt zero-shot to arbitrary branch structures with multiple forking patterns despite the presence of point cloud noise and leaf-induced clutter unseen in training. The trajectories exhibit novel characteristics, effectively using the forearm, upper and lower wrists, gripper, and the unactuated end-effector. In some instances, we observe self-reconfiguring strategies to prop up the occluding branches, while in others, the arm inserts itself between the branches and the line to separate them. Further, these results validate our RL-based approach in discovering efficient strategies and adapting to complex scenarios, well beyond what humans can demonstrate.

### v2

@@ III-G Observation & Reward Space
Proprioceptive joint pose: q{q} 6
@@
Our observation space $O$ consists of five feature sets, listed in Table I. First, we construct two feature sets from point clouds: the KME features described in Section III-F, which characterise global structure, and a local feature group, which focuses on nearby point distances. The latter includes the Euclidean distance $d$ of the end-effector to the closest $k=5$ branch points and a safety breach indicator, computed as:
@@
A robot-power line contact doesn’t compromise the validity of our approach, as the overhanging branches act as conductors already[guggenmoos2003effects]; nevertheless, a safety margin ($d_{sm}=4$cm, to account for motion and sensor uncertainties) is introduced to reduce power line wear and tear.
A novel aspect of our method is its multi-modality, attributable to a whole-arm touch detector. Recently, [jacobgentle] has shown that in the absence of expensive tactile sensors, time-series sliding window torque measurements: $\tau_{t},\tau_{t-1},\tau_{t-2},...,\tau_{t-(m-1)}]$ over the last $m$ time steps along with other proprioceptive features, can be used to build a classifier trained to detect contact ‘bumps’ at the current time-step $t$, from labeled touch data. While [jacobgentle] is trained with touch features for a simple reach task, we extend that approach and demonstrate that the touch classifier is useful even in the presence of vision and can aid complex manipulation tasks. During simulation, a proxy classifier $\mathbb{I}(\|{F_{t}}\|_{2}>f_{u})$ computes the force threshold $f_{u}$ breaches by the robot links $l$, leveraging the net contact force $F_{t}\in\mathbb{R}^{l\times 3}$ exposed by Isaac Gym. The full body contact awareness from this touch detector is crucial for the arm to leverage its multiple links (e.g. forearm and wrist) for manipulation, to supplement the partial, noisy, and time-varying point clouds.
Our simple reward structure listed below has three components: a) an occlusion clearance reward $r_{h}$, b) a smoothness reward $r_{q}$ on the joint velocities $\dot{q}$, and c) a safety bonus $r_{sm}$ to keep the arm clear of the power line safety margin,
@@
The clearance reward $r_{h}\in[0,1]$ is a monotonically decreasing function of the occlusion heuristic $h_{t}$. The $r_{sm}$ coefficient (0.4), is hand-tuned to make the sparse safety bonus noticeable without overshadowing clearance reward, whereas $r_{q}$ is a penalty term to encourage smoother trajectories.
@@
Figure 3: Terminal state poses: (a)(b): Simulation trajectory states showing the whole arm being utilised to shield the power line. (c)-(e): Similar real strategies using various arm links for clearance, executed on branches of different tree species.
@@ IV-A Simulation Setting
For power line clearance (Fig. 2b), we run 6144 parallel Gym environments, each simulating a tree with deformable branches (in green), a rigid power line (in red), and a Kinova arm actuated via velocity control of its six joints, omitting the end-effector. Each tree is formed by drawing from a Gaussian distribution ($\sigma=0.1$) representing the L-system morphology parameters to ensure high diversity in topology but sufficient physical feasibility, in addition to arbitrary trunk rotation to expose the agent to varying branch occlusion patterns. Furthermore, the power line location is randomised, but constrained to the reach region of the gripper. Finally, we perform all experiments with an average $h_{t}\in[0,1]$ of at least $0.7$ across all environments, noting that occlusion score can increase as well from the arm pushing more branches closer to the power line.
@@
A critical challenge within this framework is the contact explosion resulting from the combinatorial interactions among the multi-link arm, multi-branch tree, and the power line, exacerbated by a large number of training environments. This issue, common in contact-rich parallel simulations, can quickly overwhelm the simulation constraint solvers, triggering object inter-penetrations and CUDA memory overflows. We take a shortcut to contact reduction at the cost of a marginal drop in test performance. Specifically, during training, we allow the branches (but not the robot) to penetrate the power line through collision masking, thereby removing one key set of interactions. In contrast, during simulation tests, all contacts are in place to accurately reflect real-world, but the environment count is set low. This approach, along with careful tuning of the Temporal Gauss-Seidel (TGS) contact solver parameters, enables us to run large-scale training without compromising representational accuracy.
TABLE II: Baseline Comparison & Feature Ablation
@@
Description $d_{\text{max}}$ Train Rew Trials Test Rew Test SR% Occ Drop% Touch% Steps in Succ
(a) Baseline: Guided IK - Lift - - 1536 - 33.4 ± 2.7 12.7 ± 1.4 11.3 ± 0.4 179.1 ± 13.5
(b) Baseline: Guided IK - Sweep - - 1536 - 41.1 ± 1.7 17.2 ± 2.2 11.0 ± 0.6 193.5 ± 17.5
(c) Baseline: GNN Features 0.02 1613.9 1536 1511.7 ± 21.0 55.5 ± 2.9 33.6 ± 1.7 10.5 ± 0.1 291.4 ± 14.6
[…]
(e) (d) + Local PC Features 0.02 1531.7 1536 1486.5 ± 49.1 43.8 ± 4.5 25.2 ± 3.6 2.1 ± 0.4 258.3 ± 32.1
(f) (e) + Touch Feature 0.02 1606.6 1536 1506.1 ± 28.5 46.1 ± 2.6 27.9 ± 4.8 2.6 ± 0.8 268.9 ± 17.7
(g) (e) + KME Features 0.02 1739.1 1536 1630.2 ± 31.6 55.7 ± 1.1 40.1 ± 0.4 1.9 ± 0.1 348.2 ± 22.7
(h) All Features (RBF) 0.02 1704.3 1536 1655.9 ± 26.4 63.4 ± 2.8 46.3 ± 3.3 3.8 ± 0.3 363.4 ± 20.6
(i) All Features (Laplace) 0.02 1592.8 1536 1559.0 ± 35.0 52.5 ± 1.8 36.0 ± 2.4 4.0 ± 0.2 315.9 ± 21.0
(j) All Features (Matérn-3/2) 0.02 1662.5 1536 1620.0 ± 25.5 59.8 ± 1.3 43.5 ± 1.5 3.2 ± 0.2 362.9 ± 16.7
(k) Final Model 0.005 1825.1 1536 1763.6 ± 21.5 68.4 ± 2.1 53.7 ± 2.2 3.3 ± 0.4 429.8 ± 25.6
@@ IV-B Hardware Design & Sim-to-Real
In this work, we aimed to train a policy in simulation, running on an NVIDIA RTX 4090, and transfer it to real without a policy adaptation phase. During inference, the agent interacts with our real Kinova 6-DOF Jaco 2 through the Kinova ROS API and a custom REST interface, running on a separate low-grade workstation at a 60Hz control rate. Unlike prior works focused on artificial plants with soft leafy resistance, our test branches are randomly chosen from real trees, ensuring sufficient stiffness to resist clearance. A single fixed Intel RealSense D405 camera, operating at 30 fps, provides the RGB-D ($480\times 640$) observation of the workspace, illustrated in Fig. 3:(c)-(e). The segmented real branch point clouds are constructed from the RGB-D camera images; in contrast, simulation tree points are randomly sampled from the coarse-grained branch segment surfaces. The real robot point cloud is built once from meshes, just as in the simulation, and updated at each step with the 3D coordinate transformation tree from ROS.
Furthermore, to address the significant sim-to-real discrepancies in robot parameters, contact dynamics, and gravity compensation, exacerbated by velocity control, we leverage Segmented Steady State Error Control (SSED)[jacobgentle]. This scheme alternates between applying an action $a_{t}\sim\pi_{w}(\cdot|s_{t})$ to the desired state $s^{d}_{t}$, instead of the current, but periodically synchronising $s^{d}_{t}$ to $s_{t}$, in effect, rejecting short-term steady-state errors and clearing long-term accumulated offsets.
@@ IV-C Real Vision Pipeline
This section outlines the key transformation steps required to convert the real-time streaming images into RL features. First, we use Grounding DINO [liu2023grounding] to locate the objects in the scene with simple text prompts. The resulting bounding boxes are passed to a Segment Anything SAM-HQ [ke2023segment] model to generate object masks. The masks are then de-projected to the robot frame, leveraging the depth channel input. Crucially, manipulators operating in velocity drive mode have a control rate lower bound, 60Hz in our case, required to maintain velocity. Current state-of-the-art grounding and segmentation tools are unable to achieve this without compromising the mask quality. Therefore, we use a video segmentation framework Cutie [cheng2024putting], to propagate the mask across time-steps, occasionally replacing it with updated masks to prevent quality loss.
@@ V EXPERIMENTS AND RESULTS
Experiment 1: First, we establish quantitative non-RL baselines by executing hand-tuned, multi-step, inverse kinematics (IK) strategies in simulation. The results are shown in Table II, rows (a)-(b). In the first case (Guided IK - Lift), we drive the end-effector to a location just above the power line mid-point (where the occlusion is maximum, on average), and then lift the arm up, to push the branches away. In the next case (Guided IK - Sweep), the arm first aligns the gripper to the axis connecting the robot base and tree trunk, then executes a sweeping motion to the left, if power line is on the left of the robot, or to the right otherwise, clearing branches on the way.
As a second baseline, in Table II: (c), we leverage a GNN model [kim2024towards], where nodes represent branch fork positions and edges denote parent–child cylindrical links. However, unlike their graph-to-graph contact policy, which requires a predefined target deformation state (unknown in our context), we feed all node, edge, and global attributes to an RL policy, using only the representation component of the graph.
Experiment 2: Second, we perform a simulation feature ablation study, with RL policies, to justify our choices. For each combination, we train independent policies by injecting varying levels of Gaussian Noise $\mathcal{N}(0,\sigma^{2})$ into the point clouds with the $3\sigma$ rule, i.e., $\sigma=d_{\text{max}}/3$, $d_{\text{max}}$ being the maximum deviation representative of the camera depth inaccuracy. The feature ablation results are shown in Fig. 4 and Table II: (d)-(h). The box plot data points are policies constructed with $d_{\text{max}}$ levels ranging from $0.005$ to $0.04$, displaying the distribution of test rewards and success rates on a single unseen test set. On the other hand, each row in Table II is an individual policy tested on multiple test sets, listing a variety of additional evaluation metrics, with rows (e)-(j) showing the results applying a median noise $d_{\text{max}}=0.02$, representative of the ablation study. By contrast, row (k) shows our best model with a lower $d_{\text{max}}=0.005$, the depth inaccuracy of the D405 camera used for real tests. Furthermore, to justify our choice of RBF kernel, we ablate against two alternatives: the Laplace kernel $k_{\mathrm{L}}(x,x^{\prime})=\exp\!\left(-\|x-x^{\prime}\|/\gamma\right)$ and Matérn-3/2 kernel $k_{\mathrm{M}}(x,x^{\prime})=\left(1+\sqrt{3}\|x-x^{\prime}\|/\gamma\right)\exp\!\left(-\sqrt{3}\|x-x^{\prime}\|/\gamma\right)$, where, $\gamma>0$ is a length-scale hyper-parameter. The kernel ablation results are in Table II: (i)-(j).
@@
Experiments 1 & 2 Results: The Table II result metrics are described as follows: Applicable only to RL, Training Reward (Train Rew) and Test Reward (Test Rew) assess the agent’s performance during training and on unseen test scenarios. Test Success Rate (Test SR %) quantifies the percentage of test environments (Trials) where all occlusions were removed from the power line. Given that the branch collection will continue to resist clearance even after success is reached, we choose a conservative Test SR, considering a task as a success only if the line is occlusion-free for at least 10 consecutive steps. Occlusion reduction (Occ Drop %) is the percentage drop in heuristic ($h_{t}$) from the start to end of test episodes averaged across all environments, while the (Touch %) reflects the frequency of undesirable instances where the arm breaches the power line safety margin $\delta_{sm}$. The metric (Steps in Succ) measures the efficiency of successful trajectories; as an example, the final row (k) implies that, in a $1000$ step trajectory, the arm starting from the home position, traversed to the occlusion locations and moved the occluding branches away from the line in 1000429.8=570.21000-429.8=570.2 steps, while the remaining $429.8$ steps were completely occlusion free for potential pruning, despite continued cluster resistance. All metrics except the power line touch (Touch %) are higher, the better.
From Table II: (a)-(c), the human-specified trajectories consistently underperformed our learning strategy. Our approach also outperformed the GNN simulation baseline; moreover, the graph method is less transferable to real, as it relies on true branch node positions that must be tracked across time-steps. Predictably, both baselines resulted in significantly higher power line contacts due to the absence of a touch indicator or a contact penalty. The improved Steps in Succ for RL policies indicate that, following successful de-occlusion, the RL agent attempts to stabilise the scene by holding the branches steady, unlike guided swipes.
@@
From the feature ablation study, the plot in Fig. 4 demonstrates that both the KME feature group capturing the global point cloud attributes and the contact detection classifier form essential constituents of our multi-modal policy, and removal of either would hurt performance. Table. II: (d)-(h) demonstrates that despite the stochasticity of individual policies and high noise injection, progressive addition of feature groups induces improvement across all evaluation metrics. To highlight, we report all success rates without distinguishing between feasible and infeasible tasks, i.e., all occlusion patterns are randomly generated and contain scenarios that cannot be solved by a single arm. In some unsolvable failure instances, the arm cannot generate sufficient torques (due to joint limits) to overcome the collective branch resistance; in others, multiple disconnected branches can occlude the line at locations not closely spaced, only one of which can be removed at a time.
@@
Experiment 3: Third, we compare the representational efficiency of our approach to other point cloud feature extractors, namely, PointNet++ [qi2017pointnet++] and the Point2Vec [zeid2023point2vec]. Specifically, we replace the four global point features from Table I, i.e., $\overline{\mu_{\mathbb{P}}}[P^{wbr}]$, $\overline{\mu_{\mathbb{P}}}[P^{zbr}]$ , $\overline{\mu_{\mathbb{P}}}[P^{clr}]$ and $\overline{\mu_{\mathbb{P}}}[P^{rob}]$ with embeddings from PointNet++ and Point2Vec. In each case, we use pre-trained checkpoints, down-sample the input, and employ PCA post-feature extraction for consistency with our approach. We train each set for 250 epochs with 1024 environments, tabulating the time taken (Train Time) and the accumulated reward (Train Rew). In addition, we compute $GF_{t}^{(E)}$} denoting the average time taken to build the 4x global point features only for a single RL time-step for $E$ environments. $GF_{t}^{(1024)}$ is indicative of the representation efficiency while training on 1024 environments and $GF_{t}^{(1)}$ represents the inference performance for one environment. The results are in Table IV, where lower is better for the time metrics, such as, $GF_{t}^{(1024)}$, $GF_{t}^{(1)}$, and Train Time while higher values are better for others.
@@
Comparing the time taken for a single time-step and for the overall training, our kernel-based approach shows orders of magnitude improvement, with similar training rewards. For the PointNet++ and Point2Vec, the maximum possible environments (Max Envs) we could spin up in our hardware was 1024, an indication of the GPU memory constraint from heavy pre-trained checkpoints. Notice that our final KME-based model training, row (k) from Table II, which uses 6x environments and 5x iterations compared to this experiment, takes just 4.5 hours to complete on an RTX 4090. Furthermore, $GF_{t}^{(1)}$ values indicate that both PointNet++ and Point2Vec cannot meet the 60Hz real-time inference frequency requirement, justifying our approach. However, we acknowledge that improved efficiency may be at the cost of expressivity and that advanced baseline versions or quantisation techniques can reduce this time difference.
@@
Experiment 4: Finally, we go beyond simulation validation to evaluate our approach in a real-world laboratory setting. We test three branches from different species, modifying either the branch orientation or the power line position in each experiment. Furthermore, in each test, we start with an $h_{t}$ of at least $0.8$, and consider both partial occlusion clearance and any touch to the power line as failure instances.
@@
The results (Table IV) and the corresponding videos (supplement & Fig. 3) demonstrate that the simulation-trained policy can adapt zero-shot to arbitrary branch structures with multiple forking patterns despite the presence of point cloud noise and leaf-induced clutter unseen in training. The trajectories exhibit novel characteristics, effectively using the forearm, upper and lower wrists, gripper, and the unactuated end-effector. We observe self-reconfiguring strategies to prop up the occluding branches, while in others, the arm inserts itself between the branches and the line to separate them. Further, these results validate our RL approach in discovering efficient strategies and adapting to complex scenarios, well beyond what humans can demonstrate.